The Age Of Two Children Are 11 And 8 years. In How Many Years' Time Will The Product Of Their Age Be 208?
Answer by TimothyLamb(4379)

x = jack age now = 11
y = jill age now = 8
z = number of years from now when the product of their ages will be 208
---
(x + z)(y + z) = 208
(11 + z)(8 + z) = 208
88 + 19z + zz = 208
zz + 19z + 88 - 208 = 0
zz + 19z - 120 = 0
---
the above quadratic equation is in standard form, with a=1, b=19 and c=-120
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 19 -120
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
z = 5
z = -24
---
the negative root doesn't fit the problem statement, so use the positive root:
---
answer:
z = number of years from now when the product of their ages will be 208 = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php